An onion crate with a weight of 225 N accelerates along a frictionless surface as the crate is pulled with a force of 14.5 N as shown in the drawing. What is the horizontal acceleration of the crate?
1
Expert's answer
2020-08-28T09:33:14-0400
"m=\frac{W}{g}\\ma=F\cos{19}\\\frac{W}{g}a=F\cos{19}\\\frac{225}{9.8}a=14.5\cos{19}\\a=0.597\frac{m}{s^2}"